# Designing quantum chemistry algorithms with just-in-time compilation

### v5

@@ 1 Introduction
Traditional quantum chemistry software [16, 33, 30] has historically relied on ahead-of-time (AOT) compilation and monolithic code generation, with tightly coupled Fortran or C++ implementations [21, 8, 10] tuned for a fixed set of use cases. Although this design has served well for decades, it poses serious limitations in some modern high-performance computing (HPC) environments where GPUs are the primary source of computational power. Algorithms in electronic structure theory often require handling a combinatorial variety of tensor contraction patterns and sparsity conditions that depend on the input molecular system. Static compilation must conservatively support all conditional branches, hyperparameter combinations, and numerical thresholds, leading to bloated binaries, excessive control flow divergence, poor cache and register utilization, and inefficient use of memory bandwidth. On GPUs, these inefficiencies are further magnified by strict constraints on register pressure, warp divergence, and hardware occupancy, often requiring significant manual optimization. For example, most quantum chemistry packages deal with low-angular-momentum integrals successfully but still struggle with the performance of high-angular momentum integrals, which suffer from all of these problems.
@@
Beyond runtime performance, just-in-time compilation also offers substantial advantages to developers. JIT facilitates rapid prototyping and testing: since code is generated at runtime, one can easily benchmark different kernel variants, tiling strategies, or instruction sequences, without recompiling and linking the entire application. This accelerates development cycles, simplifies debugging, and lowers the barrier to contributing performance improvements across architectures. Furthermore, the developers are able to design novel algorithms which are not general enough under the AOT framework. In this paper, we demonstrate several algorithms due to this benefit. The development experience can be even further improved: Within a narrow area of application such as machine learning, SQL and GPU programming, one can design domain-specific languages (DSLs) via LLVM [15]. Frameworks like TensorFlow[1], PyTorch[24], JAX[5], and TVM incorporate their own DSLs—such as TensorFlow’s GraphDef and XLA HLO, PyTorch’s TorchScript and FX, JAX’s JAXpr, and TVM’s TIR—to represent algorithms in intermediate forms that enable aggressive optimization and efficient execution. These DSLs allow developers to write high-level code while benefiting from low-level performance tuning through JIT compilation and backend-specific code generation, bridging the gap between productivity and performance.
@@
JIT compilation benefits mixed precision algorithms in quantum chemistry, which naturally accelerate electron repulsion integrals by exploiting the Schwarz inequality [17, 20]. Most tasks can be calculated in single-precision (FP32) as compared to the commonly-used double-precision (FP64). Single precision algorithms benefit quantum-chemistry computations in three key ways: (1) deep learning inference GPUs and consumer-grade GPUs typically feature far more FP32 units than FP64 units, consume less power, and are more readily available; (2) single-precision data occupy half the storage, allowing more values to reside in shared memory and registers, improving data locality and throughput; and (3) hardware-accelerated FP32 implementations of the exponential (exp) and error function (erf) offer 4×–8× speedups, alleviating what is a computational bottleneck in certain workloads. These benefits can be better utilized by JIT compilation without significant changes to the code.
@@
Contributions: 1) We introduce JIT to the quantum chemistry community and optimize the two-electron integral kernels (1q1t in Sec. 3.2) with compilation techniques. Our framework integrates seamlessly with the quantum chemistry package GPU4PySCF [16, 33] and is released as the open-source JoltQC library. 2) We develop a novel fragmentation algorithm for high angular momentum integrals, improving data locality and alleviating memory-bandwidth bottlenecks through multilevel reduction. 3) We benchmark single-precision integral kernels under our JIT framework and demonstrate substantial performance gains when combining JIT compilation with our new algorithmic strategies.
@@ 2 A short formulation of Rys quadrature
$(ij|kl)=\sum_{pqrs}c_{p}c_{q}c_{r}c_{s}(pq|rs)$ (1)
$c_{p}$, $c_{q}$}, $c_{r}$, and $c_{s}$ are the pre-defined contraction coefficients for each shell in the basis set. The number of Gaussian functions in each primitive shell is $l(l+1)/2$ or $2l+1$ for cartesian atomic orbitals or spherical atomic orbitals respectively, where $l$ is the angular momentum of the shell. Only Cartesian orbitals are considered in the following sections, and spherical orbitals are just a linear combination of Cartesian orbitals with fixed transformation coefficients. These linear transformations can be performed efficiently outside the kernel. As a result, each dimension of the primitive integral $(pq|rs)$ is $N^{f}=l(l+1)/2$. We denote the dimension of the integral for each shell quartet as $(N_{i}^{f},N_{j}^{f},N_{k}^{f},N_{l}^{f})$.
@@
where $N_{r}$ is the number of Rys roots. $t_{r}$ and $w_{r}$ are the roots and weights of Rys quadrature. The roots and weights depend on the exponents and center (atomic) positions of the four primitive Gaussian functions, the total angular momentum, and a range-separation parameter $\omega$ if needed. The calculation of Rys roots and weights relies on interpolations from tabulated data, and the details have been documented in [9]. The number of Rys roots is determined at the compilation time, and only the data[33, 26] associated with the number of roots are compiled into the binary. The primitive integral at each root position can be factored into intermediate variables in three spatial directions:
@@
$I_{x}$, $I_{y}$, and $I_{z}$ are the intermediate variables introduced in [7, 9]. Each of them is a tensor of shape $(l_{i}+1,l_{j}+1,l_{k}+1,l_{l}+1)$. They depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only. They can be efficiently evaluated using recursive relations, documented in [9]. But storing those values on the limited resources on GPU can be tricky, and several strategies will be discussed in Sec. 3.1 and Sec. 3.2. The index set for enumerating the Guassian functions of a primitive shell is defined as
@@
The idx function is defined differently in various quantum chemistry packages. One will need to double check the conventions to align the quantities in atomic basis, such as density matrix, molecular orbitals, and so on. We adopt the convention of reversed lexicographic order. The indexing operation in general makes the memory access random and less efficient.
Because of its huge size, the four-dimensional integral tensor is not formed explicitly in the main storage, but instead contracted with the density matrix $D$ to form the Coulomb ($J$) and exchange ($K$) potential matrices:
@@
For a molecular system with real-valued orbitals, the 8-fold symmetry of $(ij|kl)$ can be used to save computational cost. In particular, for the integral element $(ij|kl)$ composed of the $i$-th, $j$-th, $k$-th and $l$-th atomic orbital, it is contracted with two density matrix elements $D_{kl}$ and $D_{lk}$ to form two matrix elements $J_{ij}$ and $J_{ji}$, and it is contracted with four density matrix elements $D_{jl}$, $D_{jk}$. $D_{il}$, $D_{ik}$ to form four $K$ matrix elements $K_{ik}$, $K_{il}$, $K_{jk}$, $K_{jl}$.
@@
where $\Delta D$, $\Delta J$, and $\Delta K$ are the changes of density matrix, Coulomb potential, and exchange potential between two steps of SCF iterations (i.e. incremental SCF), respectively. Since $\Delta D$ is getting smaller along the SCF iterations, the screening is more effective. In practice, the density matrix screening is performed at shell pair level, which is similar to Eq. (5).
@@ 3 Design and Algorithms
The input of a quantum chemistry calculation contains the protocol of the method, geometry of a molecule, and various thresholds. These inputs are unknown when a typical quantum chemistry package is compiled. Therefore, all types of input have to be handled as dynamic variables. The idea of JIT design is to feed the input information to the compiler at compilation time such that the compiler can aggressively optimize the code. On the other hand, if the compilation completely depends on all the input information, i.e. if we recompile for every job even if successively jobs are only different in geometry, the recompilation may slow down the overall execution time, since the compilation can be more expensive than the actual calculation with an improper design. In the design of JIT programs, distinguishing between static and dynamic variables is crucial for user experience and performance. In general, we follow this principle: The information of quantum chemistry protocol and device information are static for compilation. That is, when the quantum chemistry method, basis set, or type of GPUs is changed, the kernels should be recompiled. The same kernels will be used for different molecules. All the molecule information, including the atomic numbers, atomic types and atomic coordinates, and the total charge and spin of the system, are dynamic variables for compilation.
More specifically, we select the angular momentum and contraction pattern of the basis set, as well as the task type, as the template parameters for the kernels. They are described in Table 1. Once these variables are fixed, the size of any intermediate variable in the integral kernels is known at compilation time. If those parameters are changed, we have to recompile the kernels. With such choices, we avoid complicated runtime logic flows for different data type, task type, and unrolling strategies. Thus the amount of source code is minimal: only ~1,000 lines of code for each algorithm, much less than the ~20,000 lines in GPU4PySCF v1.4.
Template
Parameter
Data Type Description
@@
The parameters of the basis set are dynamic variables. Since many elements share the same contraction pattern but different coefficients and exponents, we do not need to compile multiple kernels for the same contraction pattern. Generally, reading those data is not the bottleneck of the kernels. All of the dynamic parameters are summarized in 2.
Input
Data Type Description
@@ 3.1 One quartet one thread (1q1t) algorithm
The first algorithm we present is largely the same as the classical Rys quadrature algorithm on CPUs [26], except for a more aggressive loop unrolling strategy is used. First, a list of significant shell quartets is generated by running over all possible combinations screened by the Schwartz inequality. Each quartet is assigned to one thread for evaluating all elements of the integral in Eq. (1) (including all primitives in a contraction and all functions in a shell), and contract it with the density matrix according to Eq. (4), until the pool of all significant shell quartet is empty. This algorithm was implemented in the early version of GPU4PySCF (v1.0 and before). This straightforward implementation has two performance issues. 1) There are still multiple loops not unrolled due to unknown variables at compilation time, such as number of primitivies for each shell and number of density matrices. This is actually the main bottleneck of low-angular-momentum integrals. 2) Significant register spill out onto local memory for high angular momentum. In this subsection, we will show that with more known variables at compilation time, the performance is significantly improved. The second issue will be discussed in the next subsection.
@@
In the single-precision implementation, the exponents, contraction coefficients, and atomic coordinates are prepared and loaded in FP32. The density matrix values are loaded as FP64 and then casted to FP32 in kernels. Inside the kernel, all the arithmetic operations are performed in FP32 and accelerated with the --use_fast_math compiler flag. One should notice that fast-math intrinsics for single precision is much faster, but not IEEE-754 accurate. When $J$ and $K$ are calculated, they are first casted to FP64, then accumulated with atomicAdd in FP64.
We benchmark the performance of JK kernels with a straight peptide chain composed of 120 glycines (gly120), and 6-31G [6] basis set. Density matrix screening is turned off for profiling purpose. The timings on A100-80G and A10-24G are shown in Figure 2 and 3 respectively. The overall speedup of FP64 against AOH compilation in GPU4PySCF v1.4 is about 2.0x. For low-angular-momentum, the evaluations of exp and erf functions are often the computational bottleneck. Thanks to the special hardware acceleration of these functions on GPU, the overall speedup of FP32 reaches about 5-10x on A100-80G. Since A10-24G has 1/32 hardware units for FP64 compared to FP32, FP64 algorithms are always bounded by the limited compute units. The improvement with JIT is marginal. When switched to FP32, the 1q1t algorithm is accelerated by 30x.

### v6

@@ 1 Introduction
Traditional quantum chemistry software [17, 34, 31] has historically relied on ahead-of-time (AOT) compilation and monolithic code generation, with tightly coupled Fortran or C++ implementations [22, 8, 10] tuned for a fixed set of use cases. Although this design has served well for decades, it poses serious limitations in some modern high-performance computing (HPC) environments where GPUs are the primary source of computational power. Algorithms in electronic structure theory often require handling a combinatorial variety of tensor contraction patterns and sparsity conditions that depend on the input molecular system. Static compilation must conservatively support all conditional branches, parameter combinations, and numerical thresholds, leading to bloated binaries, excessive control flow divergence, poor cache and register utilization, and inefficient use of memory bandwidth. On GPUs, these inefficiencies are further magnified by strict constraints on register pressure, warp divergence, and hardware occupancy, often requiring significant manual optimization. For example, most quantum chemistry packages deal with low-angular-momentum integrals successfully but still struggle with the performance of high-angular-momentum integrals, which suffer from all of these problems.
@@
Beyond runtime performance, just-in-time compilation also offers substantial advantages to developers. JIT facilitates rapid prototyping and testing: since code is generated at runtime, one can easily benchmark different kernel variants, tiling strategies, or instruction sequences, without recompiling and linking the entire application. This accelerates development cycles, simplifies debugging, and lowers the barrier to contributing performance improvements across architectures. Furthermore, developers can explore specialized algorithms that would be impractical to hand-code under the AOT model. In this paper, we demonstrate several algorithms due to this benefit. The development experience can be even further improved: Within a narrow area of application such as machine learning, SQL and GPU programming, one can design domain-specific languages (DSLs) via LLVM [16]. Frameworks like TensorFlow[1], PyTorch[25], JAX[5], and TVM incorporate their own DSLs—such as TensorFlow’s GraphDef and XLA HLO, PyTorch’s TorchScript and FX, JAX’s JAXpr, and TVM’s TIR—to represent algorithms in intermediate forms that enable aggressive optimization and efficient execution. These DSLs allow developers to write high-level code while benefiting from low-level performance tuning through JIT compilation and backend-specific code generation, bridging the gap between productivity and performance.
@@
The Schwarz inequality provides rigorous upper bounds on integral magnitudes, enabling most shell quartets to be evaluated in single precision (FP32) rather than the conventional double precision (FP64) [18, 21]. JIT compilation makes this mixed-precision approach particularly easy to implement: since precision is a compile-time parameter, the same source code generates both FP32 and FP64 kernels without code duplication. Most integral tasks can be calculated in single-precision as compared to the commonly-used double-precision. Single precision algorithms benefit quantum-chemistry computations in three key ways: (1) deep learning inference GPUs and consumer-grade GPUs typically feature far more FP32 units than FP64 units, consume less power, and are more readily available; (2) single-precision data occupy half the storage, allowing more values to reside in shared memory and registers, improving data locality and throughput; and (3) hardware-accelerated FP32 implementations of the exponential (exp) and error function (erf) offer 4×–8× speedups, alleviating what is a computational bottleneck in certain workloads. JIT compilation exploits all three advantages automatically: the same kernel source is compiled once for FP32 and once for FP64, with the compiler selecting the appropriate hardware intrinsics and register allocation for each precision — no manual code duplication required.
@@
Contributions: 1) We introduce JIT to the quantum chemistry community and optimize the two-electron integral kernels (1q1t in Sec. 3.1 and 1qnt in Sec. 3.2) with compilation techniques. Our framework integrates seamlessly with the quantum chemistry package GPU4PySCF [17, 34] and is released as the open-source JoltQC library. 2) We develop a novel fragmentation algorithm for high angular momentum integrals, improving data locality and alleviating memory-bandwidth bottlenecks through multilevel reduction. 3) We benchmark single-precision integral kernels under our JIT framework and demonstrate substantial performance gains when combining JIT compilation with our new algorithmic strategies.
@@ 2 A short formulation of Rys quadrature
$(ij|kl)=\sum_{p\in i}\sum_{q\in j}\sum_{r\in k}\sum_{s\in l}c_{ip}\,c_{jq}\,c_{kr}\,c_{ls}\;(pq|rs).$ (1)
$c_{ip}$, $c_{jq}$, $c_{kr}$, and $c_{ls}$ are the pre-defined contraction coefficients for each shell in the basis set. The number of Gaussian functions in each primitive shell is $(l+1)(l+2)/2$ or $2l+1$ for cartesian atomic orbitals or spherical atomic orbitals respectively, where $l$ is the angular momentum of the shell. Only Cartesian orbitals are considered in the following sections, and spherical orbitals are just linear combinations of Cartesian orbitals with fixed transformation coefficients. These linear transformations can be performed efficiently outside the kernel. As a result, each dimension of the primitive integral $(pq|rs)$ is $N^{f}=(l+1)(l+2)/2$. We denote the dimension of the integral for each shell quartet as $(N_{i}^{f},N_{j}^{f},N_{k}^{f},N_{l}^{f})$.
@@
where $N_{\text{r}}=\left\lfloor(l_{i}+l_{j}+l_{k}+l_{l})/2\right\rfloor+1$ is the number of Rys roots. $t_{r}$ and $w_{r}$ are the roots and weights of Rys quadrature. The roots and weights depend on the exponents and center (atomic) positions of the four primitive Gaussian functions, the total angular momentum, and a range-separation parameter $\omega$ if needed. The calculation of Rys roots and weights relies on interpolations from tabulated data, and the details have been documented in [9]. The number of Rys roots is determined at the compilation time, and only the data [34, 27] associated with the number of roots are compiled into the binary. The primitive integral at each root position can be factored into intermediate variables in three spatial directions:
@@
$I_{x}$, $I_{y}$, and $I_{z}$ are the intermediate variables introduced in [7, 9]. For every root, each of the intermediate variables is a tensor of shape $(l_{i}+1,l_{j}+1,l_{k}+1,l_{l}+1)$. The total number of intermediate variables is $3\times 3\times(2\times 2\times 2\times 2)=144$ for calculating a $(pp|pp)$ integral (the first 3 refers to the three Cartesian dimensions, and the second 3 is the number of roots). For $(gg|gg)$ integral, the number of variables is raised up to $3\times 9\times(5\times 5\times 5\times 5)=16875$, which exceeds the capacity of GPU registers. The size of intermediate variables can be reduced if the roots are handled sequentially. The same space can be reused for different roots. The intermediate variables $I_{x},I_{y},\text{and }I_{z}$ depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only. They can be efficiently evaluated using recursive relations, documented in [9]. But storing those values on the limited resources on GPU can be tricky, and several strategies will be discussed in Sec. 3.1 and Sec. 3.2. The index set for enumerating the Gaussian functions of a primitive shell is defined as
@@
The idx function is defined differently in various quantum chemistry packages. We adopt reversed lexicographic order: for $l=2$ (d shell), $\texttt{idx}(2,0,0)=0$, $\texttt{idx}(1,1,0)=1$, …, $\texttt{idx}(0,0,2)=5$. Note that other packages may adopt different orderings; the density matrix and orbital coefficients must be reordered accordingly. The indexing operation in general makes the memory access random and less efficient.
Because of its huge size, the four-dimensional integral tensor is not formed explicitly in the main memory, but instead contracted with the density matrix $D$ to form the Coulomb ($J$) and exchange ($K$) potential matrices:
@@
For a molecular system with real-valued orbitals, the 8-fold symmetry of $(ij|kl)$ can be used to save computational cost,
(ij|kl)=(ji|kl)=(ij|lk)=(ji|lk)=(kl|ij)=(lk|ij)=(kl|ji)=(lk|ji).(ij|kl)=(ji|kl)=(ij|lk)=(ji|lk)=(kl|ij)=(lk|ij)=(kl|ji)=(lk|ji). (5)
As a result, only integrals $(ij|kl)$ where $i>j$, $k>l$ and $iN_{ao}+j>kN_{ao}+l$ needs to be evaluated, and this saves $7/8$ of the integral evaluation work. In particular, for the integral element $(ij|kl)$ composed of the $i$-th, $j$-th, $k$-th and $l$-th atomic orbital, the J matrix requires a trace over the bra-ket pair $(k,l)$, summing $(ij|kl)D_{kl}$, yielding two updates ($J_{ij}$ and $J_{ji}$ by symmetry). The K matrix involves cross-contractions where one index from the bra pairs and one from the ket, yielding four distinct updates. Therefore, the integral $(ij|kl)$ is contracted with two density matrix elements $D_{kl}$ and $D_{lk}$ to form two matrix elements $J_{ij}$ and $J_{ji}$. And it is simultaneously contracted with four density matrix elements $D_{jl}$, $D_{jk}$. $D_{il}$, $D_{ik}$ to form four $K$ matrix elements $K_{ik}$, $K_{il}$, $K_{jk}$, $K_{jl}$.
@@
where $\Delta D$, $\Delta J$, and $\Delta K$ are the changes of density matrix, Coulomb potential, and exchange potential between two steps of SCF iterations (i.e. incremental SCF), respectively. Since $\Delta D$ decreases along the SCF iterations, the screening is more effective. In practice, the density matrix screening is performed at shell pair level, which is similar to Eq. (6).
@@ 3 Design and Algorithms
The central design decision in a JIT framework is which inputs to treat as compile-time constants (static) and which to pass as runtime arguments (dynamic). Making too few variables static forfeits optimization opportunities; making too many static causes excessive recompilation. We adopt the following rule: protocol and device information — the quantum chemistry method, basis set type, and GPU architecture — are static, while molecular information — atomic coordinates, Gaussian orbital exponents, and contraction coefficients — remains dynamic. This ensures that a single compilation serves all molecules sharing the same basis set, while still enabling aggressive loop unrolling and register allocation.
More specifically, we select the angular momentum and contraction pattern of the basis set, as well as the task type, as the template parameters for the kernels. They are described in Table 1. A crucial consequence of this design: once the angular momentum, contraction pattern, and number of primitives are fixed, every loop bound and array size in the integral kernel becomes a compile-time constant. The compiler can then fully unroll loops, allocate registers statically, and eliminate all branching. If those parameters are changed, we have to recompile the kernels. With such choices, we avoid complicated runtime logic flows for different data types, task types, and unrolling strategies. Thus the amount of source code is minimal: only ~1,000 lines of code for each algorithm, much less than the ~20,000 lines in GPU4PySCF v1.4.
Template Parameter Data Type Description
@@
The parameters of the basis set are dynamic variables. Since many elements share the same contraction pattern but different coefficients and exponents, we do not need to compile multiple kernels for the same contraction pattern. The basis set parameters (exponents, coefficients, coordinates) are stored in global memory and read once per shell quartet. Since the data size is usually small and is efficiently cached through the L1/L2 cache, reading those data is not the bottleneck of the kernels. Another strategy is to form the basis pair data as the kernel input, and the pair data can be more efficient in exp operation in some scenarios, for example F64 kernel on RTX cards, at the cost of less efficient cache utilization. Directly taking the basis data as input is part of the current design, but may not be optimal.
All of the dynamic parameters are summarized in 2.
Input Data Type Description
@@
Based on this design, we build heuristics (including algorithms and fragmentation strategies) to deliver optimized kernels for a specific device type and precision type (FP64/FP32). The heuristics are benchmarked ahead of time, and the best ones are saved in a database as a part of the final package. In the runtime, the kernel caller will find the corresponding algorithm based on the device type and required precision. Then, only one optimized kernel is compiled. It should be noted that, even for the exact same integrals, with different device/precision combination, different heuristics might be selected. In particular, the FP64 algorithm may be different from the FP32 algorithm for the same angular momentum, in order to achieve the best performance.
In the next two sections, we present two heuristics algorithms for integral kernels, each targeting a different angular-momentum regime. The 1q1t algorithm (Sec. 3.1) assigns one thread per shell quartet and is often well-suited for low angular momentum (s, p shells), where the integrals fit in registers. The 1qnt algorithm (Sec. 3.2) fragments the integral across multiple threads and is designed for high angular momentum (d, f, g shells), where register pressure becomes the bottleneck.
@@ 3.1 One quartet one thread (1q1t) algorithm
The first algorithm we present is largely the same as the classical Rys quadrature algorithm on CPUs [27], except that a more aggressive loop unrolling strategy is used. First, a list of significant shell quartets is generated by running over all possible combinations screened by the Schwartz inequality. Each quartet is assigned to one thread for evaluating all elements of the integral in Eq. (1) (including all primitives in a contraction and all functions in a shell), and contract it with the density matrix according to Eq. (4). This algorithm was implemented in the early version of GPU4PySCF (v1.0 and before). This straightforward implementation has two performance issues. 1) Several performance-critical loops — over primitives in each shell and over density matrices — have bounds unknown at compile time and therefore cannot be unrolled. On GPUs, unrolled loops enable the compiler to schedule instructions more efficiently, eliminate branch overhead, and allocate registers to specific loop iterations. This is the dominant bottleneck for low-angular-momentum integrals, which are individually cheap but numerous. JIT compilation resolves this by making all loop bounds compile-time constants. 2) Significant register spills to local memory for high angular momentum. In this subsection, we show that with more known variables at compilation time, the performance is significantly improved. The second issue will be discussed in the next subsection.
@@
In the single-precision implementation, the exponents, contraction coefficients, and atomic coordinates are prepared and loaded in FP32. The density matrix is maintained in FP64 throughout the SCF procedure for diagonalization and DIIS, so we load it in its native format and cast to FP32 inside the kernel to avoid maintaining a separate single-precision copy. One may improve the memory throughput by maintaining a separate single-precision copy. Inside the kernel, all the arithmetic operations are performed in FP32 and accelerated with the --use_fast_math compiler flag. The --use_fast_math flag replaces standard math functions (sqrt, exp, erf) with hardware-accelerated intrinsics that are faster but not IEEE-754 compliant (i.e. they introduce an additional small rounding error). The accelerated exp function benefits the evaluations of Gaussian functions, as the kernel takes the native Gaussian exponents and contraction coefficients as inputs. The pair data is calculated inside the kernel. The benefit of exp is negligible if one precompute the pair data before invoking the JK kernel. The calculations of Rys roots and weights are accelerated as well for two reasons. 1) We directly use erf to evaluate the root and weight when $N_{r}=1$ (interpolation is used for other cases), as we found that directly evaluating the special functions is often faster than the interpolation. 2) In the long-distance regime for $N_{r}>1$ (outside of interpolation window) of Rys roots and weights, sqrt is used in the asymptotic formula. [13]
We benchmark the performance of JK kernels with a straight peptide chain composed of 120 glycines (gly120), and 6-31G [6] basis set. Density matrix screening is turned off for profiling purpose. The timings on A100-80G and A10-24G are shown in Figure 2 and 3 respectively. The overall speedup of FP64 against AOH compilation in GPU4PySCF v1.4 is about 2.0x. For low-angular-momentum, the evaluations of exp and erf functions are often the computational bottleneck. Thanks to the special hardware acceleration of these functions on GPU, the overall speedup of FP32 reaches about 5-10x on A100-80G. Since the A10-24G only delivers 1/32 of P64 throughput compared to FP32, FP64 algorithms are always bounded by the limited compute units. The improvement with JIT is marginal. When switched to FP32, the 1q1t algorithm is accelerated by 30x.